Find two 2 x 2 upper triangular matrices C,D such that C and D are both
not diagonal, but CD is diagonal.
1
Expert's answer
2020-01-30T11:53:43-0500

For example:

"C=\begin{bmatrix} 1 &1 \\0&1
\end{bmatrix}, \;
D=\begin{bmatrix} 1 &-1 \\0 &1
\end{bmatrix},
\; 
C D=\begin{bmatrix} 1 &0\\0 &1
\end{bmatrix}"


You can find other examples:


"C=\begin{bmatrix} a &b \\0&c
\end{bmatrix}, \;
D=\begin{bmatrix} d &-b \\0 &a
\end{bmatrix}
, \; 
CD=\begin{bmatrix} ad &0 \\0&ac
\end{bmatrix}" , where a,b,c,d≠0